OpenStudy (anonymous):

Differential Equations power series. y''+xy'-3y=0 x_(0) = 0 obtain up to the 13th term of the power series... i need help getting through this problem.. i have worked on it for 5 hours today and i seem to not get the right answer.

OpenStudy (anonymous):

I am sorry for not understanding it right away, but you need a power series representation of a family of solutions or Piccardi's method or?

OpenStudy (anonymous):

i haven't seen any of that in my text,but it talks about Airy's equation. http://ltcconline.net/greenl/courses/204/powerlaplace/seriessolutions1.htm i have found this example online and its similar to my problem but i can seem to get a answer i am satisfied with..

OpenStudy (anonymous):

Ok, I will give it a shot then. I think I understand what you need :-)

OpenStudy (eyust707):

\[ y = \sum_{0}^{\infty} c _{n} x^n \]

OpenStudy (anonymous):

Ok see i have these and this is what you need for these problems. \[y = \sum_{n=0}^{\infty} A _{n}x ^{n}\] \[y ' = \sum_{n=1}^{\infty} nA _{n}x ^{n-1}\] \[y'' = \sum_{n=2}^{\infty}n(n-1)A _{n}x ^{n-2}\]

OpenStudy (eyust707):

yes

OpenStudy (eyust707):

you plug those in

OpenStudy (eyust707):

then set your x power all to the same one

OpenStudy (anonymous):

i have these then i need to apply these but where i am getting stuck is when im trying to get my power the same.

OpenStudy (anonymous):

I think you can get to this: \[\large 2a_2 - 3a_0 \sum_{n=1}^{\infty}[(n+2)(n+1)a_{n+2} + na_n - 3a_n]x^n = 0\]

OpenStudy (eyust707):

so the y' one becomes x^n cause there is an x alreayd there

OpenStudy (anonymous):

Typo, there should be a + between the a0 and the sum :-)

OpenStudy (anonymous):

From there, I think you have to use the inital condition, though.

OpenStudy (eyust707):

nope you use the initial condiditon at the end

OpenStudy (eyust707):

you need to plug in values of n

OpenStudy (eyust707):

wait are you sure its x(0)=0 and not y(0) = 0?

OpenStudy (anonymous):

yes i have reached that far with that same equation.. but i find myself solving a3 a1 and a3 and such.. and thats where im really confused

OpenStudy (anonymous):

\[x _{0} =0\]

OpenStudy (eyust707):

oh

OpenStudy (anonymous):

From there, I think that's the way @eyust707 said, plug in values for n, and try to find a general form for a series representation. But I will leave it to you, guys. I need to go to sleep. I will check back if the problem remains unsolved.

OpenStudy (anonymous):

ok thanks man.

OpenStudy (anonymous):

No problem. Wasn't too helpful tho :-)

OpenStudy (anonymous):

well at least i am not totally off.

OpenStudy (eyust707):

I can explain how to solve this but its very difficult to do by typing..

OpenStudy (eyust707):

i bet theres a patrickjmt video on almost exactly the same thing

OpenStudy (anonymous):

ok on youtube?

OpenStudy (eyust707):

http://www.youtube.com/watch?v=6csP7dw0XTY

OpenStudy (anonymous):

ok thanks man i will try to figure this out now. and i be back on tomorrow if i need anymore help! thanks much!

OpenStudy (eyust707):

if you are cant get it tommorrow and you really want to understand it add me on skype and ill explain it to you. my name is the same on there

OpenStudy (anonymous):

ok man i will i cant remember mine right now, but if you get someone random it will probably be me. Thanks

OpenStudy (anonymous):

thanks guys i think i got it...